OpenStudy (aravindg):

prove : (p-1)(q-r)+(q-1)(r-p)+(r-1)(p-q) looking for both algebraic and logical explanations

OpenStudy (aravindg):

=0

OpenStudy (aravindg):

i mean the above expression equal to 0

OpenStudy (aravindg):

prove

OpenStudy (anonymous):

are q,p and r in ap?

OpenStudy (aravindg):

??

OpenStudy (aravindg):

no

OpenStudy (anonymous):

i won't be able to prove it until and unless you give me a relationship between p,q and r

OpenStudy (aravindg):

i need a logical explanation

OpenStudy (anonymous):

Okay hmm let me try

OpenStudy (aravindg):

hey we can

OpenStudy (anonymous):

aravind did you even open it

OpenStudy (mr.math):

\[(p-1)(q-r)+(q-1)(r-p)+(r-1)(p-q)\] \[=pq-rp-q+r+qr-pq-r+p+rp-rq-p+q=0.\]

OpenStudy (mr.math):

Just cancel out like-terms with different signs.

OpenStudy (anonymous):

just open it you will get zero lol

OpenStudy (aravindg):

i knw that u needed a logical explanation involving property of cyclic